7 The publisher of a sports magazine plans to offer new subscribers one of three gifts: a sweatshirt with the logo of their favorite team, a coffee cup with the logo of their favorite team, or a pair of earrings also with the logo of their favorite team. In a sample of 431 new subscribers, the number selecting each gift is reported below. At the 0.02 significance level, is there a preference for the gifts or should we conclude that the gifts are equally well liked? 12.5 points Gift Sweatshirt Coffee cup Earrings Frequency 197 123 Skipped eBook References Нол, r- Hi The proportions are not equal. State the decision rule. Use 0.01 significance level. (Round your answer to 3 decimal places.) Reject HO if chi-square> How many degrees of freedom are there? Degrees of freedomCompute the value of chi-square. (Negative amounts should be indicated by a minus sign. Leave no cells blank- be certain to enter O wherever required. Round なand fo-fe to 2 decimal places and (fo-ね2/g to 4 decimal places. Round X to 4 decimal places.) 12.5 points Skipped Gift Sweatshit Coffee Cup Earrings Total fo fe fo fe (fo - fe)A2/fe eBook 431 431 What is your decision regarding Ho? IS There a preference for the gifts.
Answer #1

a) REJECT H0 if chi square> 9.210 at alpha=0.01

degrees of freedom= n-1=3-1=2

FREQUENCY fo fe fo-fe (fo-fe)2 (fo-fe)A2/fe GIFT sweatshirt Coffee cup Earrings TOTAL 197 143.67 53.33 2844.089 123 143.6720.67 427.2489 111 143.67 -32.67 1067.329 431 431.01 19.796 2.9738 7.429 30.1988

HO IS REJECTED THERE IS A PREFERENCE FOR THE GIFTS.